# Time Evidence Fusion Network: Multi-source View in Long-Term Time Series Forecasting

### v3

@@ 4.8 Robustness Experiments
TEFN is a highly robust neural network. To verify the robustness of TEFN in real prediction, a comparative experiment was conducted by adding noise to the dataset. The unit noise $\gamma$ follows the Gaussian distribution $N(0,1)$ with an expectation of 00 and a variance of $1$. Assume that the input time series is $x$, with time and channel dimensions. To adapt to the scale of the data, the transformation as Eq.(18) is applied to input data where $i$ is the index of channel dimension, $i$ is the index of time dimension, $x_{i}$ is a vector containing all time steps of channel $i$ and $std(\cdot)$ is Standard deviation function. So, the noisy data experiments are compared with the original data experiments.
$x^{\prime}_{ij}=x_{ij}+std(x_{i})*\gamma$ (18)
@@
The experimental results of robustness testing are shown in the Table V. Continuing from the analysis, the experimental outcomes underscore the robust nature of TEFN, where the introduction of Gaussian noise to the dataset led to only marginal increases in prediction errors, with most variations being insignificant relative to the original errors. Notably, in long-term forecasting scenarios, a counterintuitive improvement in accuracy was observed in some instances, indicating that TEFN’s adaptability and learning capacity in complex, dynamic environments extend beyond mere resilience to noise, potentially leveraging subtle perturbations to enhance prediction performance.
TABLE V: The results of robustness experiments
@@
Weather 96 0.182 0.227 0.232 ($\uparrow$ 0.050) 0.269 ($\uparrow$ 0.042)
@@
4.9 Interpretability Analysis
The Time Evidence Fusion Network (TEFN) incorporates fuzzy logic and the Dempster-Shafer theory to effectively model the inherent uncertainty and ambiguity within time series data, resulting in a model that is not only accurate but also highly interpretable. TEFN allocates mass to events within the event space $E=\{E_{i}|i\in[1,2^{|S|}]\}$, where each event represents a potential outcome or scenario. The mass assigned to an event reflects the level of confidence or support for that event, providing insight into the model’s belief about its likelihood of occurrence. This support allocation is particularly interpretable due to the use of fuzzy membership functions $\mu(x)=wx+b$ within the BPA module, as depicted in Figure 9. By analyzing the shape and parameters of these membership functions, users can gain a clear understanding of the model’s perception of data features and their influence on predictions. For instance, it becomes evident which features are deemed more important than others, and how the model responds to variations within these features. This visualization not only makes the model’s inner workings more transparent but also allows users to intuitively assess the model’s confidence in its predictions and the level of uncertainty associated with each outcome.
@@ 4.9 Interpretability Analysis
4.10 Limitation and Future Prospects
TEFN is a naive and unmodified model. There is not engineering experience in the application of BPA modules, such as how to generate BPA, and fuzzy logic is not the optimal implementation method. Due to fuzzy logic, TEFN is essentially a linear model that cannot handle some nonlinear time series well. In addition, for the fusion of mass distribution in the face of the failure of the classic DSR for a large amount of data, using expectations as the fusion method may not necessarily be the optimal pattern. In terms of performance optimization, TEFN adopts a simple fully connected approach, which can lead to a rapid increase in parameter count for time series that are too long or have multiple channels. In fact, the implementation techniques of TEFN can further optimize the performance of the model through sampling and kernel operations similar to convolution. Thanks to the fact that evidence theory belongs to symbolic logic, the interpretation of neural network parameters will be more in line with human logic. Effective parameter initialization may exist, for example, initializing corresponding parameters in a non-neural network way may lead to faster convergence.
5 Conclusion

### v4

@@ IV-H Robustness Experiments
TEFN is a highly robust neural network. To verify the robustness of TEFN in real prediction, a comparative experiment was conducted by adding noise to the dataset. The unit noise $\varepsilon$ follows the Gaussian distribution $N(0,1)$ with an expectation of 0 and a variance of $1$. Assume that the input time series is $x$, with time and channel dimensions. To adapt to the scale of the data, the transformation as Equation (19) is applied to input data where $i$ is the index of channel dimension, $i$ is the index of time dimension, $x_{i}$ is a vector containing all time steps of channel $i$ and $std(\cdot)$ is Standard deviation function. So, the noisy data experiments are compared with the original data experiments.
$x^{\prime}_{ij}=x_{ij}+\text{std}(x_{i})*\varepsilon$ (19)
@@
The experimental results of robustness testing are shown in the Table VI. Continuing from the analysis, the experimental outcomes underscore the robust nature of TEFN, where the introduction of Gaussian noise to the dataset led to only marginal increases in prediction errors, with most variations being insignificant relative to the original errors. Notably, in long-term forecasting scenarios, a counterintuitive improvement in accuracy was observed in some instances, indicating that TEFN’s adaptability and learning capacity in complex, dynamic environments extend beyond mere resilience to noise, potentially leveraging subtle perturbations to enhance prediction performance.
TABLE VI: The results of robustness experiments
@@
Weather 96 0.182 0.227 0.232 ($\uparrow$ 0.05) 0.269 ($\uparrow$ 0.042)
@@
IV-I Interpretability Analysis
(a) Channel BPA
(b) Time BPA
The Time Evidence Fusion Network (TEFN) incorporates fuzzy logic and the Dempster-Shafer theory to effectively model the inherent uncertainty and ambiguity within time series data, resulting in a model that is not only accurate but also highly interpretable. TEFN allocates mass to events within the event space $E=\{E_{i}|i\in[1,2^{|S|}]\}$, where each event represents a potential outcome or scenario. The mass assigned to an event reflects the level of confidence or support for that event, providing insight into the model’s belief about its likelihood of occurrence. This support allocation is particularly interpretable due to the use of fuzzy membership functions $\mu(x)=wx+b$ within the BPA module, as depicted in Figure 10. By analyzing the shape and parameters of these membership functions, users can gain a clear understanding of the model’s perception of data features and their influence on predictions. For instance, it becomes clear which features are deemed more important than others, and how the model responds to variations within these features. This visualization not only makes the model’s inner workings more transparent but also allows users to intuitively assess the model’s confidence in its predictions and the level of uncertainty associated with each outcome.
@@ IV-I Interpretability Analysis
IV-J Limitation
TEFN is a naive and unmodified model. There is not engineering experience in the application of BPA modules, such as how to generate BPA, and fuzzy logic is not the optimal implementation method. Due to fuzzy logic, TEFN is essentially a linear model that cannot handle some nonlinear time series well. The experiment shows that adding a nonlinear activation function can effectively improve the experimental performance of TEFN. In addition, for the fusion of mass distribution in the face of the failure of the classic DSR for a large amount of data, using expectations as the fusion method may not necessarily be the optimal pattern. In terms of performance optimization, TEFN adopts a simple fully connected approach, which can lead to a rapid increase in parameter count for time series that are too long or have multiple channels. In fact, the implementation techniques of TEFN can further optimize the performance of the model through sampling and kernel operations similar to convolution. Thanks to the fact that evidence theory belongs to symbolic logic, the interpretation of neural network parameters will be more in line with human logic. Effective parameter initialization may exist, for example, initializing corresponding parameters in a non-neural network way may lead to faster convergence.
V Conclusion